# Enhancing operational wind downscaling capabilities over Canada: Application of a Conditional Wasserstein GAN methodology

### v1

@@ 5 Conclusions
In this study, we presented a downscaling methodology based on a conditional Wasserstein GAN (WGAN-GP), leveraging low-resolution data from the GDPS and high-resolution data from the HRDPS, conditioned on high-resolution static covariates. We also conducted experiments incorporating the Frequency Separation approach to enhance the performance of the model. To implement the conditional WGAN-GP, we employed a UNET model as the generator. This UNET was designed to process two types of inputs: high-resolution static covariates and low-resolution GDPS data. Due to memory constraints, the training protocol relied on extracting random crops from single pairs of GDPS and HRDPS data. We carried out experiments to analyze the impact of including high-resolution static covariates in the model. Our findings indicate that these covariates significantly improved model convergence. Furthermore, experimental results demonstrated that the use of Frequency Separation contributed to further performance gains. We evaluated the generalization capability of the proposed model using a held-out test set. The results revealed that the best model configuration incorporated Frequency Separation, achieving favorable performance in terms of RMSE and Log Spectral Distance (LSD) metrics. To operationalize the model, we proposed a downscaling procedure for the entire Canadian domain. This procedure generates overlapping downscaled patches of size $128\times 128$ and blends them into the larger Canadian region using a linear interpolation approach. As future work, we aim to operationalize the model as a product capable of extending high-resolution wind forecast services from the current 48-hour horizon to the 10-day forecast horizon of the Canadian global forecast model. Additionally, we plan to investigate the application of recent state-of-the-art weather foundation models [37] for wind downscaling.

### v3

@@ 5 Conclusions
In this study, we presented a downscaling methodology based on a conditional Wasserstein GAN (WGAN-GP), leveraging low-resolution data from the GDPS and high-resolution data from the HRDPS, conditioned on high-resolution static covariates. We also conducted experiments incorporating the Frequency Separation approach to enhance the performance of the model. To implement the conditional WGAN-GP, we employed a UNET model as the generator. This UNET was designed to process two types of inputs: high-resolution static covariates and low-resolution GDPS data. Due to memory constraints, the training protocol relied on extracting random crops from single pairs of GDPS and HRDPS data. We carried out experiments to analyze the impact of including high-resolution static covariates in the model. Our findings indicate that these covariates significantly improved model convergence. Furthermore, experimental results demonstrated that the use of Frequency Separation contributed to further performance gains. We evaluated the generalization capability of the proposed model using a held-out test set. The results revealed that the best model configuration incorporated Frequency Separation, achieving favorable performance in terms of RMSE and Log Spectral Distance (LSD) metrics.
In preparation for an operational environment, we trained the AI model on small $128\times 128$ image patches using a downsampling factor of 8 and made inferences over the entire Canadian domain $1288\times 2536$. As future work, we aim to develop the model as a product capable of extending high-resolution wind forecast services from the current 48-hour horizon to the 10-day forecast horizon of the Canadian global forecast model. Additionally, we plan to investigate the application of recent state-of-the-art weather foundation models [43] for wind downscaling.